OpenStudy (anonymous):

Which function has zeros at x = 10 and x = 2? f(x) = x2 – 12x + 20 f(x) = x2 – 20x + 12 f(x) = 5x2 + 40x + 60 f(x) = 5x2 + 60x + 100

OpenStudy (anonymous):

A function g(x) has x-intercepts at (1/2, 0) and (6, 0). Which could be g(x)? g(x) = 2(x + 1)(x + 6) g(x) = (x – 6)(2x – 1) g(x) = 2(x – 2)(x – 6) g(x) = (x + 6)(x + 2)

OpenStudy (susanna):

(x-10)(x-2) x^2-12x+20

OpenStudy (anonymous):

thank you so much do you know the other one ?

OpenStudy (susanna):

I'm working on that...

OpenStudy (susanna):

It's g(x) = (x – 6)(2x – 1)

OpenStudy (anonymous):

thank you so much for you help omg

OpenStudy (nincompoop):

I hope you are learning a lot with this cheating going on

OpenStudy (anonymous):

im retaking a class it dont matter lol